Assume that when adults with smartphones are randomly selected, 44% use them in meetings or classes. If 7 adult smartphone users are randomly selected, find the probability that at least 3 of them use their smartphones in meetings or classes.


 The probability is _______ Round to four decimal places as needed.)  
Answer #1

Solution :

Given that p = 0.44 , n = 7

=> q = 1 - p = 0.56

=> P(x >= 3) = 1 - P(x < 3)

= 1 - [P(x = 0) + P(x = 1) + P(x = 2)]

= 1 - [7C0*(0.44^0)*(0.56^7) + 7C1*(0.44^1)*(0.56^6) + 7C2*(0.44^2)*(0.56^5)]

= 1 - [0.0173 + 0.0950 + 0.2239]

= 1 - 0.3362

= 0.6638

=> The required probability is 0.6638